OpenStudy (anonymous):

how to solve √6x+5=√53

OpenStudy (anonymous):

Perhaps it should be written as: \[\sqrt{6x+5}=\sqrt{53}\] squaring both sides we find. 6x+5= 53 6x=53-5 6x= 48 x= 48/6 x= 8

OpenStudy (anonymous):

@Nisha12

OpenStudy (anonymous):

thnxxxz... :) u from which class??

OpenStudy (anonymous):

@Nisha12 Where r u from?